find the volume of a frustum of a cone with bases of radii 5cm and 19cm. The height of this frustum is 6cm.: find the volume of a frustum of a cone with bases of radii 5cm and 19cm. The height of this frustum is 6cm.
Answer by orca(336)

Let h be the height, R the radius of the lower base, and r the radius of the upper base,then its volume is:
V = (pi*h/3)(R^2+R*r+r^2)
Substituting
R = 19
r = 5
h = 6
into the formula, we have
V = (pi*h/3)(R^2+R*r+r^2)
= (3.14*6/3)(19^2+19*5+5^2)
= 3032.68(cm)^3